### v1

# CORE: Lossless Compression for Retrieval-Augmented LLMs via Reinforcement Learning

Abstract
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance the timeliness of knowledge and the factual accuracy of responses in Large Language Models (LLMs). However, the inclusion of excessive retrieved documents substantially increases the input length, leading to higher computational costs. Previous studies have attempted to compress retrieved documents into shorter texts before in-context integration, but such methods often compromise end-task performance. The lack of well-defined compression targets forces many approaches to rely on fixed heuristics, which cannot guarantee that the compressed content will effectively support the end task. To address these limitations, we propose CORE, a novel method designed to achieve lossless context compression for RAG. CORE employs reinforcement learning to optimize the compression process without relying on predefined compression labels. Specifically, it utilizes end-task performance as a reward signal and applies Generalized Reinforcement Learning Policy Optimization (GRPO) to train the compressor. This end-to-end training framework enables the compressor to generate summaries that maximize the accuracy of answers generated by the LLM. Extensive experiments on four datasets demonstrate the superiority of our approach. With a high compression ratio of 3%, our method not only avoids performance degradation compared to prepending full documents across all datasets but also improves the average Exact Match (EM) score by 3.3 points. The code will be released soon.
1 Introduction
Large language models (LLMs) have undergone rapid development in recent years, significantly enhancing performance across various language tasks due to their emergent capabilities in semantic understanding and reasoning. Nevertheless, LLMs still face challenges in updating knowledge and providing factual responses Fan et al. (2024). To address these issues, Retrieval-Augmented Generation (RAG) has emerged as a promising approach Ram et al. (2023). By retrieving the most relevant documents from extensive knowledge bases and prepending them as contextual information to the original input, RAG substantially improves LLM performance on knowledge-intensive tasks Ram et al. (2023); Fan et al. (2024).
As shown in Figure 1, in-context retrieval-augmented generation significantly outperforms the baseline without RAG (i.e., the number of docs equals to zero). As the number of prepended retrieved documents increases, the accuracy continues to rise, eventually saturating when the number of documents reaches 10. However, this improvement in accuracy introduces two limitations: a substantial rise in computational costs due to the increased number of tokens that must be encoded, and challenges for the model in effectively utilizing all information in the context, often resulting in the omission of key information located in the middle Liu et al. (2023).
To mitigate these issues, recent research has made strides in compressing retrieved documents Xu et al. ; Cao et al. (2024); Jin et al. (2024b); Zhu et al. (2024); Kim & Thorne (2025); Wu et al. (2025); Jin et al. (2024a); Zhang et al. (2024a). These methods mainly include summarizing retrieved documents Xu et al. , extracting key information Xu et al. ; Cao et al. (2024), constructing key supporting evidence Jin et al. (2024b), and filtering noise based on information theory Zhu et al. (2024). Although these efforts have shown promising progress, they still face several limitations. Firstly, most methods achieve document compression at the expense of performance degradation; for instance, RECOMP Xu et al. exhibits a decline of 3–5 exact match (EM) points after compression (Table 1), making it unsuitable for accuracy-sensitive scenarios. Secondly, despite serving downstream tasks, the absence of ideal compression targets prevents most existing methods from being optimized in an end-to-end manner. This limitation hinders the compressor’s ability to generate content that reliably supports the LLM in producing accurate responses, thereby undermining end-task performance. Finally, some approaches employ compressors with parameter counts comparable to those of the target LLM itself, resulting in non-negligible computational costs that may offset the efficiency gains of compression.
To overcome these limitations, we propose COmpression via REinforcement learning (CORE), a novel method designed to achieve lossless context compression for retrieval-augmented generation. Recently, reinforcement learning (RL) has emerged as a promising approach for enhancing reasoning capabilities without requiring supervised data on intermediate reasoning steps Liu et al. (2024); Chen et al. (2025a). Inspired by this advancement, we employ RL to train a compressor that has significantly fewer parameters than the main LLM to summarize documents. Specifically, the summaries generated by the compressor are concatenated with the original input and fed into the LLM to obtain predicted answers. A rule based reward signal is then formulated based on answer accuracy, i.e., end task performance, to optimize the compressor. The training process is implemented using the Group Policy Optimization (GRPO) Shao et al. (2024) algorithm, which effectively integrates the reward signal derived from end-task performance to update the compressor in a stable and efficient manner. This end-to-end design enables the compressor to be trained in a goal-oriented manner even in the absence of ideal compression labels, ensuring that the generated summaries effectively assist the LLM in producing accurate responses.
We evaluate the effectiveness of CORE on four benchmark datasets: two single-hop question-answering datasets, Natural Questions and TriviaQA, along with two multi-hop datasets, HotpotQA and 2WikiMultihopQA. Experimental results (Table 1) demonstrate that CORE achieves state-of-the-art performance compared to all baselines. While maintaining a high compression ratio of 3%, our method not only avoids performance degradation relative to prepending full documents across all four datasets, but also improves the average Exact Match (EM) score by 3.3 points. Furthermore, we empirically show that our approach is not dependent on a specific compressor architecture (Figure 5). The trained compressor also exhibits strong transferability to other large language models (Table 2). Finally, via a detailed case study (Table 4 and 5), we highlight the advantages of our compression method over alternative approaches.
Figure 1: Growth of EM and F1 scores with increasing retrieved documents on two datasets.
2 CORE-RAG
Figure 2: Overview of our method.
In this section, we introduce our proposed method, CORE-RAG. First, we provide an overview of the entire workflow. Then, we detail how to train a compressor using reinforcement learning to significantly reduce the number of document tokens while maintaining performance.
2.1 Pipeline
We adopt the same problem formulation with prior work Xu et al. , where given an input sequence $x$, target output sequence $y$, and a set of $N$ retrieved documents $D$, our objective is to compress $D$ with respect to $x$ into a summary $s$ that preserves the most relevant information while using significantly fewer tokens than $D$. The whole process consists of two key components: a compression model $c_{\theta}\colon(x,D)\mapsto s$ and a large language model $M\colon(x,s)\mapsto y$ that generates the output sequence $y$. We treat $M$ as a black-box system while focusing exclusively on training the compressor. The compressor employed in this study is also a language model. However, it is intentionally designed to be significantly smaller than $M$, as we aim to reduce computational costs of encoding the retrieved documents.
2.2 Training the Compressor with RL
Our compressor is designed to generate document summaries while ensuring that these summaries are particularly useful for the base LLM to perform a downstream task. The latter requirement is extremely challenging as the criteria for an effective summary is inherently task-dependent, meaning that no supervised training data is available to guide the compression. To address this challenge, we employ reinforcement learning, motivated by its demonstrated success in training language models without explicit intermediate supervision. The overview of our training method is shown in Figure 2. The remainder of this section elaborates on our RL training methodology, including distillation warm-up, policy optimization, and reward formulation.
2.2.1 Distillation for RL Warm-Start
Since our chosen compression model has a relatively small number of parameters, its query-focused document summarization capability is limited. To provide reinforcement learning with a strong initial policy and ensure training stability, we first employ knowledge distillation from a teacher model to initialize the RL policy. Specifically, we use an extremely large language model (DeepSeek-V3) as the teacher model to generate summaries of the retrieved document sets in response to the given query, which serve as target summaries. We then evaluate the end task performance under two conditions: (1) with the target summary prepended to the input, and (2) with the original input alone (i.e., no retrieval). By comparing these results on the base model ($M$), we filter out instances where prepending the summary does not improve performance. The remaining samples are used for supervised fine-tuning.
2.2.2 Group Relative Policy Optimization
After the warm-up phase, we employ reinforcement learning to train the compressor in an end-to-end manner. Our objective is to optimize the summaries generated by the compressor to improve the performance of the end task. However, obtaining perfect summary labels is challenging, as even summaries generated by extremely large language models cannot be guaranteed to be the most effective. Therefore, we leverage reinforcement learning to optimize the compressor, allowing it to learn how to generate summaries that enhance end-task performance.
Specifically, we employ GRPO Shao et al. (2024) as the learning algorithm. Unlike Proximal Policy Optimization (PPO), which trains a separate critic model, GRPO estimates the baseline directly from a group of rollouts. Given an existing policy, $\pi_{\theta_{\text{old}}}$, and a reference policy, $\pi_{\theta_{\text{ref}}}$, the GRPO objective maximizes the policy $\pi_{\theta}$ using $G$ rollouts $\tau=\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)$, for each input $x\sim\mathcal{D}$:
$\mathcal{J}(\theta)$ $={\mathbb{E}}_{x\sim\mathcal{D},\ \{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}$ (1)
$\frac{1}{G}\sum_{i=1}^{G}\Bigg{[}\min\Bigg{(}\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}A_{i},\ \text{clip}\Bigg{(}\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)},1-\epsilon,1+\epsilon\Bigg{)}A_{i}\Bigg{)}-\beta\mathbb{D}_{\text{KL}}\big{(}\pi_{\theta}\parallel\pi_{\theta_{\text{ref}}}\big{)}\Bigg{]},$
where $A_{i}=(r_{i}-\text{mean}(\{r_{j}\}_{j=1}^{G}))/\text{std}(\{r_{j}\}_{j=1}^{G})$ represents the normalized advantage of the $i$-th rollout within the group, $\epsilon$ is the clipping ratio, and $\beta$ is the coefficient for the KL divergence penalty. The inclusion of the KL divergence term ensures that the updated policy does not deviate significantly from the reference policy.
2.2.3 Reward
  • Generating End-Task Output. It is important to note that the reward is not calculated directly based on the output summary of the compressor. Instead, the summary generated by the compressor is prepended to the original input (question) and then fed into the base LLM $M$ to produce an output $y$. The reward is subsequently calculated based on $y$. Throughout the entire training process, $M$ is fixed and does not participate in training.
  • Computing Rewards. We design simple rule-based rewards based on end-task performance to guide the compressor’s improvement, which consists of two components:
    • EM Reward ($r_{\text{EM}}$). We employ exact match (EM) as a main reward function, which is a widely adopted metric for evaluating the accuracy of generated outputs. The EM reward is binary: it yields a value of 1 if the generated answer perfectly matches the ground truth, and 0 otherwise.
    • F1 Reward ($r_{\text{F1}}$). Since exact matches occur infrequently in practice, relying solely on EM-based rewards leads to sparse reward signals. Furthermore, the EM metric fails to distinguish between partially correct answers, as all non-exact matches receive zero reward. Therefore, we introduce F1 score, which provides finer-grained evaluation by measuring the degree of partial match.
    The final reward function combines these signals through weighted summation:
    $r=r_{\text{EM}}+\alpha\cdot r_{\text{F1}},$ (2)
    where $\alpha\in(0,1]$ is a hyperparameter that controls the relative contribution of the F1 metric.
3 Experiments
3.1 Experimental Settings
  • Datasets. We evaluate our method on four benchmark datasets: two single-hop question-answering datasets, Natural Questions Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017), and two multi-hop question-answering datasets, HotpotQA Yang et al. (2018) and 2WikiMultihopQA Ho et al. (2020). Results are reported on the test sets of Natural Questions and TriviaQA, as well as the development sets of HotpotQA and 2WikiMultihopQA. Performance is measured using Exact Match and token-level F1 scores for answer strings.
  • Base Large Language Models. We use Qwen2.5-14B-Instruct as the primary LLM model $M$ to generate predicted answers which are used to guide the training of the compressor. To evaluate the generalization ability of our method, we also transfer to another LLM model, LLama3.1-8B-Instruct.
  • Compressor Models. We trained our compressor model using Qwen2.5-1.5B-Instruct to generate summaries of the retrieved documents. To evaluate the effect of using different models for compression, we also trained compressors based on Llama3.2-1B-Instruct and Llama3.2-3B-Instruct.
  • Retrieval Corpus and Retrievers. Following previous studies Xu et al. , we use the Wikipedia corpus from December 20, 2018, as the retrieval source for all four datasets. The articles are segmented into non-overlapping 100-word documents. To ensure that our method is not dependent on a specific retriever, we experiment with several mainstream retrieval models. Specifically, we use DPR Karpukhin et al. (2020) for the Natural Questions dataset, a hybrid of DPR and BM25 Robertson et al. (1995) for the TriviaQA dataset, and the Contriever model Izacard et al. (2021) trained on the MS MARCO dataset Nguyen et al. (2016) for HotpotQA and 2WikiMultihopQA.
    Figure 3: Prompt template used to generate the summary.
    Figure 4: Prompt template for LLM QA.
  • Prompt Format. The prompt for the compressor is shown in Figure 3, which is used to generate summary of documents. For the question-answering prompt provided to the LLM, we include few-shot in-context examples in the prompt, followed by the retrieved documents and the question, following Xu et al. . Three in-context examples are randomly sampled from the training set. To remain consistent with Xu et al. , the retrieved documents are concatenated in ascending order of their retrieval scores, with the highest-scoring document placed closest to the question Si et al. (2022). This prompt template is provided in Figure 4.
  • Baselines. To evaluate the effectiveness of our method, we compared it against various baselines. First, we evaluated the uncompressed approach—retaining the original in-context RALM setup—by prepending the top 1, 3, 5, and 10 retrieved documents to the prompt. We also tested alternative compression methods, including the traditional BM25 algorithm (which ranks sentences by their similarity to the input), DeepSeek-V3 model (671B parameters, far exceeding our compressor’s capacity), and the off-the-shelf Qwen2.5-1.5B-Instruct model (with comparable parameter size to our approach). Additionally, we included state-of-the-art context compression methods for RAG, RECOMP Xu et al. and NoiseFilter-IB Zhu et al. (2024). For RECOMP, we evaluate both its abstractive (summary-based) and extractive (sentence selection) variants.
  • Implementation Details. The reinforcement learning framework in our study is based on verl  [^1] https://github.com/volcengine/verl. We train the compression model for two epochs on each dataset. Training is conducted on eight NVIDIA H20 GPUs using full parameter optimization, with a learning rate of 1e-5, a batch size of 256, five rollouts, and a KL loss coefficient of 0.001. For the base LLM used to generate rewards, we deploy it using vLLM. For the distillation warm-up phase, we perform full-parameter supervised fine-tuning on the compression model for two epochs using LLaMA-Factory  [^2] https://github.com/hiyouga/LLaMA-Factory.
Table 1: Open-domain QA results using Qwen2.5-14B-Instruct as the LLM. The reported token counts represent the length of in-context evidence documents, excluding few-shot examples. RECOMP, NoiseFilter-IB and our proposed CORE are all trained using Qwen2.5-1.5B-Instruct.
NQ TriviaQA HotpotQA 2WikiMultihopQA
EM F1 # tok EM F1 # tok EM F1 # tok EM F1 # tok
No Retrieval 21.36 30.97 0 53.23 59.98 0 21.05 29.48 0 26.11 29.51 0
RALM without compression
Top1 Document 34.46 44.41 142 60.82 68.70 143 29.20 38.93 147 26.79 31.87 153
Top3 Documents 37.78 48.45 427 62.60 71.02 430 31.64 41.87 442 27.89 33.58 460
Top5 Documents 38.03 49.16 712 64.10 72.48 715 32.99 43.69 737 29.64 35.21 766
Top10 Documents 38.67 50.03 1425 64.40 72.92 1428 33.95 44.88 1471 31.04 36.75 1531
Compression of top 5 documents
BM25 25.23 36.47 37 55.36 63.90 39 24.18 35.73 71 25.42 30.29 68
Qwen2.5-1.5B-Instruct 31.94 43.03 36 57.99 66.70 30 27.36 37.47 33 25.93 31.18 32
RECOMP-Abs (1.5B) 34.18 46.26 58 60.31 68.50 53 28.96 39.95 56 30.25 36.73 52
RECOMP-Ext (1.5B) 33.84 46.05 56 60.18 68.39 48 29.93 41.09 45 30.78 37.07 51
NoiseFilter-IB (1.5B) 35.15 45.94 48 59.51 68.15 35 27.97 38.62 38 27.85 34.69 40
DeepSeek-V3 (671B) 37.73 50.39 54 64.13 73.20 50 33.59 44.83 48 27.99 32.67 92
CORE (1.5B) 41.02 50.40 46 65.63 72.55 32 33.67 45.06 36 36.72 42.05 49
Compression of top 10 documents (with the compressor trained on top 5 docs)
BM25 25.91 36.88 38 55.28 63.16 37 23.49 35.01 68 25.61 30.54 65
Qwen2.5-1.5B-Instruct 32.94 44.84 40 58.45 67.31 33 28.17 38.48 36 26.22 31.57 34
RECOMP-Abs (1.5B) 34.40 46.93 59 61.42 69.88 52 31.54 42.92 52 31.98 38.16 49
RECOMP-Ext (1.5B) 33.96 46.34 60 61.03 69.51 50 31.92 43.18 55 32.52 38.87 44
NoiseFilter-IB (1.5B) 35.36 46.24 50 59.92 68.32 38 28.21 38.83 38 28.63 35.16 42
DeepSeek-V3 (671B) 37.79 51.07 56 65.29 74.45 53 34.62 45.69 50 29.00 34.64 40
CORE (1.5B) 41.88 51.26 52 66.76 73.64 33 34.68 46.35 37 37.99 43.28 48
3.2 Main Results
The detailed comparison results are presented in Table 1. Each experiment was repeated five times, and the average results are reported. We compared the following approaches: a baseline RAG method without compression using the top 1, 3, 5, and 10 documents prepended to the original input; a traditional retrieval method (BM25) for compression; off-the-shelf LLMs (Qwen2.5-1.5B-Instruct and DeepSeek-V3 (671B)); and state-of-the-art trained compressors for RAG (RECOMP-Abstractive, RECOMP-Extractive, and NoiseFilter-IB). For a fair comparison, RECOMP-Abstractive, RECOMP-Extractive, NoiseFilter-IB, and our method CORE were all trained based on the same model, Qwen2.5-1.5B-Instruct. All of these trained compressors were trained using five documents, and we report their performance both on five-document compression and their generalization to ten-document compression.
As shown in the table, prepending retrieved documents yields a substantial improvement over using no retrieval, indicating that the retrieved documents contain valuable information relevant to the answer. Performance improves as the number of retrieved documents increases, albeit with diminishing returns, which can be attributed to the decreasing relevance of lower-ranked documents. When compressing the top five documents, all baseline compression methods result in performance degradation to varying degrees compared to the uncompressed baseline. Specifically, the traditional BM25 approach leads to a significant drop in performance—for instance, a decrease of 13 EM points on NQ. Using the off-the-shelf Qwen2.5-1.5B-Instruct model for compression improves results over BM25 but still underperforms relative to the uncompressed baseline. Among the trained compressors, both RECOMP and NoiseFilter-IB exhibit a decline of 3–5 EM points compared to the uncompressed setting. We also observe that when using the extremely large-scale DeepSeek-V3 model (with 671B parameters) for compression, near-lossless compression is achieved on TriviaQA and HotpotQA; however, performance on NQ and 2Wiki remains below that of the uncompressed baseline. In contrast, our method—with only 1.5B parameters—achieves lossless compression across all four datasets and even outperforms the uncompressed baseline by 1 to 7 EM points. Overall, our approach not only significantly outperforms compression methods of comparable model size, but also maintains an advantage over the hundreds-of-times-larger DeepSeek model, demonstrating that it achieves high compression ratios without sacrificing performance—and even enhancing it.
Moreover, when generalizing the trained compressors to top-10 documents, the aforementioned conclusions remain valid. Our method continues to perform best. On NQ, it achieves a compression ratio of 3.6% in token usage while improving Exact Match (EM) by 3.2 points compared to prepending all ten documents. Similar trends are observed on TQA, with a compression ratio of 2.3% and a gain of 2.4 EM points relative to using the full set of ten documents. On HotpotQA, the method attains a compression ratio of 2.5% and an improvement of 0.7 EM points. On 2WikiMultihopQA, it achieves a compression ratio of 3.1% along with a notable increase of 6.9 EM points over prepending all ten documents.
3.3 Generalization Ability Analysis
Figure 5: The impact of different models used to train the compressor.
  • Model-Agnostic Compression Framework. Figure 5 presents a comparison of compressors trained using different models across the four datasets, while the baseline LLM used to generate the final answer remains fixed as Qwen2.5-14B-Instruct. The compressor models evaluated include LLama3.2-1B-Instruct, Qwen2.5-1.5B-Instruct, and LLama3.2-3B-Instruct, which differ in both model architecture and number of parameters. The results demonstrate that: (1) Compressors trained with our method using any of these three models achieve lossless compression and outperform the uncompressed baseline, indicating that our framework is not dependent on a specific compressor architecture and attesting to the robustness of our approach. (2) As the size of the compressor model increases, so does the compression performance, reflecting the scaling law.
    Table 2: Transfer results to Llama-3.1-8B-Instruct.
    NQ TriviaQA HotpotQA 2WikiMultihopQA
    EM F1 # tok EM F1 # tok EM F1 # tok EM F1 # tok
    No Retrieval 24.04 34.91 0 55.64 62.57 0 19.93 27.75 0 27.64 31.18 0
    RALM without compression
    Top1 Document 33.80 44.06 142 59.17 67.50 143 27.95 37.49 147 28.41 33.43 153
    Top3 Documents 36.87 47.81 427 61.13 70.06 430 30.17 40.71 442 28.67 34.23 460
    Top5 Documents 37.65 48.87 712 62.26 71.04 715 31.44 42.16 737 29.43 35.18 766
    Top10 Documents 38.12 49.93 1425 63.95 72.71 1428 32.19 42.62 1471 30.45 36.04 1531
    Compression of top 5 documents
    Qwen2.5-1.5B-Instruct 32.60 44.21 36 56.76 65.77 30 26.86 36.90 33 25.45 30.88 32
    RECOMP-Abs (1.5B) 33.41 45.50 58 58.50 67.37 53 28.85 39.76 56 31.63 37.81 52
    RECOMP-Ext (1.5B) 33.12 45.06 60 57.98 66.84 55 29.03 40.04 52 31.85 38.02 55
    DeepSeek-V3 (671B) 37.56 50.11 54 62.52 72.34 50 33.05 44.25 48 28.64 33.87 92
    CORE (1.5B) 40.72 50.00 46 64.08 71.13 32 32.17 43.71 36 35.99 41.42 49
    Compression of top 10 documents
    Qwen2.5-1.5B-Instruct 32.88 44.66 40 57.44 66.56 33 27.31 37.31 36 25.80 31.30 34
    RECOMP-Abs (1.5B) 34.18 46.80 59 59.69 68.89 52 30.17 41.42 55 33.61 39.78 44
    RECOMP-Ext (1.5B) 34.06 46.55 60 59.33 68.71 50 30.52 41.98 55 33.52 39.42 44
    DeepSeek-V3 (671B) 37.49 51.28 56 63.79 73.80 53 34.24 45.35 50 31.45 37.09 40
    CORE (1.5B) 41.77 51.27 52 65.25 72.45 33 33.25 45.09 37 37.59 42.87 48
  • Training-Free Adaptation to New LLMs. We report the results of transferring our compressor to the LLaMA-3.1-8B model in Table 2, using exactly the same input (including in-context examples and prepended document summaries) as employed with the Qwen2.5-14B-Instruct model. Note that the compressor was trained using feedback generated by Qwen2.5-14B-Instruct. The results indicate that the trained compressor baselines (e.g., RECOMP) exhibit limited generalization ability, as reflected in their increased performance gap relative to the baseline of prepending full documents. In contrast, our method demonstrates strong generalization, outperforming the baseline of prepending full documents across all four datasets while maintaining a high compression rate. This suggests that the summaries generated by our trained compressor are inherently high-quality and contain key information relevant to the answer, enabling effective transfer to other LLMs to support accurate answering.
3.4 Further Analysis
  • Table 3: Ablation study.
    Dataset Metric w/o distillation w/o GRPO CoRe NQ EM 36.37 34.18 41.02 F1 46.91 46.26 50.40 TQA EM 65.23 60.31 65.63 F1 72.41 68.50 72.55 HotpotQA EM 32.01 28.96 33.67 F1 42.73 39.95 45.06 2Wiki EM 31.40 30.25 36.72 F1 36.89 36.73 42.05
    Figure 6: The impact of $\alpha$.
  • The Impact of Distillation and GRPO. Table 3 presents an ablation study on the two key components of our method: distillation and GRPO. Here, “w/o distillation” denotes training the compressor with GRPO directly, bypassing the warm-start phase, while “w/o GRPO” indicates using only the distillation step without subsequent reinforcement learning. The results demonstrate that removing either component leads to performance degradation, confirming the necessity of both. The decline is more substantial when GRPO is omitted, highlighting the crucial role of reinforcement learning in the absence of explicit supervision. Distillation injects external knowledge into the model, providing a favorable starting point for RL training and thereby enabling RL to more effectively unlock the model’s full potential.
  • The Impact of $\alpha$. Figure 6 illustrates the performance of our method under different values of $\alpha$, where $\alpha$ controls the weighting coefficient of the F1 reward term. Setting $\alpha=0$ corresponds to using only the Exact Match (EM) reward. Results show that performance improves when $\alpha>0$ across all datasets, indicating the effectiveness of the F1 reward in alleviating the sparsity issue of the EM reward. Although the optimal value of $\alpha$ varies across datasets, values between 0.2 and 0.5 generally yield strong performance.
Table 4: Case study on NQ dataset.
Question: when did the us stop drafting for the vietnam war? Gold answer: [1973]
Top-5 documents: last men conscripted, who were born in 1952 and who reported for duty in June 1973. On February 2, 1972, a drawing was held to determine draft priority numbers for men born in 1953, but in early 1973 it was announced by Secretary of Defense Melvin Laird that no further draft orders would be issued. In March 1973, 1974, and 1975, the Selective Service assigned draft priority numbers for all men born in 1954, 1955, and 1956, in case the draft was extended, but it never was. Command Sergeant Major Jeff Mellinger, believed to be the last drafted enlisted ranked. The Gates Commission issued its report in February 1970, describing how adequate military strength could be maintained without having conscription. The existing draft law was expiring at the end of June 1971, but the Department of Defense and Nixon administration decided the draft needed to continue for at least some time. In February 1971, the administration requested of Congress a two-year extension of the draft, to June 1973. Senatorial opponents of the war wanted to reduce this to a one-year extension, or eliminate the draft altogether, or tie the draft renewal to a timetable for troop withdrawal from Vietnam; Senator. draft operated from 1940 until 1946 when further inductions were suspended, and its legislative authorization expired without further extension by Congress in 1947. During this time, more than 10 million men had been inducted into military service. However, the Selective Service System remained intact. Scattered opposition was encountered especially in the northern cities where some African-Americans protested the system. The tiny Nation of Islam was at the forefront, with many Black Muslims jailed for refusing the draft, and their leader Elijah Muhammed was sentenced to federal prison for 5 years for inciting draft resistance. Organized draft resistance also developed in. was the first time any form of student deferment was used. During the Korean War a student carrying at least twelve semester hours was spared until the end of his current semester. The United States breathed easier with the Korean War Armistice on July 27, 1953; however, technology brought new promises and threats. U.S. air and nuclear power fueled the Eisenhower doctrine of ”massive retaliation”. This strategy demanded more machines and fewer foot soldiers, so the draft slipped to the back burner. However, the head of the SSS, Maj. Gen. Hershey, urged caution fearing the conflict looming in Vietnam. In. policy of using the draft as force to compel ”voluntary” enlistment was unique in U.S. history. Previous drafts had not aimed at encouraging individuals to sign up in order to gain preferential placement or less dangerous postings. However, the incremental buildup of Vietnam without a clear threat to the country bolstered this. Some estimates suggest conscription encompassed almost one-third of all eligible men during the period of 1965–69. This group represented those without exemption or resources to avoid military service. During the active combat phase, the possibility of avoiding combat by selecting their service and military specialty led as many.
Compressor Generated summary Predicted answer
Off-the-shell Qwen2.5-1.5B-Instruct The US stopped drafting for the Vietnam War in late 1972 after the draft expired at the end of June 1971. late 1972
RECOMP The U.S. stopped drafting for the Vietnam War after the Selective Service System was officially abolished in December 1972 due to growing public opposition and the expiration of the draft’s legal authority in 1971. While the draft was formally ended, some individuals were still consripted into the military December 1972
Ours The U.S. stopped drafting for the Vietnam War in 1973 after announcing the decision by Secretary of Defense Melvin Laird earlier that year. Although the Selective Service System was later updated to include draft priority numbers, this decision marked the end of the draft’s use for national service. 1973
Table 5: Case study on 2Wiki dataset.
Question: Who is Charles Bretagne Marie De La Trémoille’s paternal grandfather? Gold answer: [Charles Armand René de La Trémoille]
Top-5 documents: as at Versailles: he was brigadier of cavalry (January 1709), first gentleman of the King’s chamber (June 1709), governor of Thouars (July 1709), and Maréchal de camp (February 1719). His sister Marie Armande Victoire de La Trémoille married Emmanuel Théodose de La Tour d’Auvergne. On 13 April 1706 he married Marie-Madeleine Motier de La Fayette (1691–1717), the daughter of Rene-Armand, marquis de La Fayette and Marie-Madeleine de Marillac, and granddaughter of the author Marie-Madeleine Pioche de la Vergne, comtesse de la Fayette. They had one child, Charles Armand René de La Trémoille, born in 1708. Charles Louis Bretagne de La Charles Bretagne Marie de La Trémoille Charles Bretagne Marie de La Trémoille (24 March 1764 – 10 November 1839), 8th duc de Thouars, was a French soldier and the son of Jean Bretagne Charles de La Trémoille and his wife, Marie-Maximilienne, princess of Salm-Kyrburg. La Trémoille married Louise-Emmanuelle de Châtillon in 1781. She was a grand daughter of Louis César de La Baume Le Blanc, the famous writer. The couple had one daughter: At the outbreak of the French Revolution, he sided with King Louis XVI, as did the rest of his family. In 1789, La Trémoille and his parents Jean Bretagne Charles de La Trémoille Jean-Bretagne-Charles de La Trémoille (5 February 1737 – 19 May 1792), 7th duc de Thouars, was a French soldier and the son of Charles Armand René de La Trémoille and his wife, Marie Hortense de La Tour d’Auvergne. La Trémoille began his military career in the Seven Years’ War as the head of the regiment of Aquitaine-Cavalry. He was seriously wounded and afterwards promoted to the rank of sergeant, and later of brigadier, of the armies of the king. During the French Revolution, La Trémoille remained faithful to King Louis XVI. He emigrated in Charles Louis Bretagne de La Trémoille Charles Louis Bretagne de La Trémoille (15 March 1683 – 9 October 1719), 6th duke of Thouars, was the son of Charles Belgique Hollande de La Trémoille and Madeleine de Créquy, daughter and heiress of Charles III de Créquy. He became duke of Thouars upon his father’s death in 1709. He was also count of Laval and Montfort and inherited the ancient French claim to the Kingdom of Naples. The Neapolitan claim gave the family the rank of ”princes étrangers” at the French court. The Duke made a career in the military as well de Saint-Didier. They had two daughters: Following her death, La Trémoille married for a third time in 1830 to Valentine Eugénie Joséphine Walsh de Serrant. They had two children: La Trémoille died shortly after his son’s birth, in 1839. Charles Bretagne Marie de La Trémoille Charles Bretagne Marie de La Trémoille (24 March 1764 – 10 November 1839), 8th duc de Thouars, was a French soldier and the son of Jean Bretagne Charles de La Trémoille and his wife, Marie-Maximilienne, princess of Salm-Kyrburg. La Trémoille married Louise-Emmanuelle de Châtillon in 1781. She was a grand daughter of Louis César de
Compressor Generated summary Predicted answer
Off-the-shell Qwen2.5-1.5B-Instruct Charles Bretagne Marie de La Trémoille’s paternal grandfather was Jean Bretagne Charles de La Trémoille. Jean Bretagne Charles de La Trémoille
RECOMP Charles Bretagne Marie de La Trémoille’s paternal grandfather was Charles Louis Bretagne de La Trémoille, who was the son of Charles Belgique Hollande de La Trémoille and Madeleine de Créquy. This makes him the direct paternal grandfather of Charles Bretagne Marie de La Trémoille. Charles Louis Bretagne de La Trémoille
Ours Charles Bretagne Marie de La Trémoille’s paternal grandfather is Charles Armand René de La Trémoille, her father’s father, the 7th duc de Thouars. Charles Armand René de La Trémoille
3.5 Case Study
To conduct an in-depth analysis of the advantages of our compressor, we performed case studies on one single-hop QA dataset (NQ) and one multi-hop QA dataset (2Wiki), with the results presented in Table 4 and Table 5, respectively. For each case, we compared the summaries generated by off-the-shelf Qwen2.5-1.5B, RECOMP, and our method CORE based on the same set of documents, as well as the final predictions made by the LLM after prepending these summaries.
As shown in the tables, although the summaries produced by off-the-shelf Qwen2.5-1.5B are concise, they largely fail to capture key information relevant to answering the question. In contrast, RECOMP demonstrates better summarization capability but is prone to being overwhelmed by lengthy documents, resulting in misjudgments and even generating misleading information—such as the statement in Table 4: “The U.S. stopped drafting for the Vietnam War after the Selective Service System was officially abolished in December 1972”—which leads the LLM to produce the incorrect answer “1972”. Our method, CORE, accurately extracts answer-critical information from lengthy documents, exemplified by the summary: “The U.S. stopped drafting for the Vietnam War in 1973 after announcing the decision by Secretary of Defense Melvin Laird earlier that year,” thereby enabling the LLM to generate the correct answer “1973”. This indicates that our compressor, trained with an end-task target-oriented reinforcement learning optimization strategy, can produce document summaries that are most helpful for answering the given question while effectively filtering out irrelevant information.
4 Related Work
4.1 Compression in RAG
RAG enhances the performance of Large Language Models (LLMs) on knowledge-intensive tasks by retrieving the most relevant documents from extensive knowledge bases and prepending them as contextual information to the original input Ram et al. (2023); Fan et al. (2024). However, this approach requires the LLM to process significantly longer token sequences, resulting in increased computational costs. To mitigate this issue, researchers have begun to explore methods for compressing retrieved documents in retrieval-augmented generation systems Xu et al. ; Cao et al. (2024); Jin et al. (2024b); Zhu et al. (2024); Kim & Thorne (2025); Wu et al. (2025); Jin et al. (2024a); Zhang et al. (2024a). Xu et al. propose compressing retrieved documents into a textual summary prior to in-context augmentation, training the compressor via data selection and distillation. Cao et al. (2024) introduce a Query-Guided Compressor (QGC), which leverages queries to guide the context compression process, thereby effectively preserving key information. Jin et al. (2024b) refine retrieved documents into Key Supporting Evidence (KSE) through knowledge synthesis, supervised fine-tuning (SFT), and preference alignment. Zhu et al. (2024) propose an information-theoretic approach that filters noise by simultaneously maximizing the mutual information between the compressed content and the ground-truth output. Kim & Thorne (2025) train a compressor to extract critical information using reward functions based on heuristic rules. However, most of these methods compromise end-task performance and lack end-to-end optimization, thus failing to ensure that the compressor maximally benefits the LLM.
4.2 Reinforcement Learning
Reinforcement learning has recently achieved significant success, enabling LLMs to develop reasoning capabilities without explicit step-by-step supervision, relying instead on rule-based reward signals Liu et al. (2024); Shao et al. (2024). Inspired by these advances, several studies have leveraged reinforcement learning (RL) to enhance retrieval-augmented generation (RAG). For instance, Kulkarni et al. (2024) employ RL to autonomously decide whether to retrieve documents or not. Zhang et al. (2024b) use RL to optimize the ranking of retrieved documents. Mao et al. (2024) propose a framework for training query rewriting models without human annotations for RAG. Chen et al. (2025b) improve retrieval-augmented generation through multi-agent reinforcement learning, incorporating a Query Rewriter, Retriever, and Generator. Additionally, a line of research applies RL to integrate search with reasoning in a step-by-step manner. Chen et al. (2025a) introduce a novel framework named ReSearch, which trains LLMs to reason with search via reinforcement learning without relying on supervised data for reasoning steps. Similar approaches include R1-Searcher Song et al. (2025), WebThinker Li et al. (2025), and DeepResearcher Zheng et al. (2025). In contrast to these works, our approach utilizes RL to train a document compressor in an end-to-end fashion, enhancing the efficiency of RAG while preserving its performance.